Section 29.7 of the text derives/explains what hapThe textbook is Physics for science and engineers by Knight 3rd edition.
Answer #1

If half of the space fills with another dielectric then its capacitance would be
= kCo / 2
and remaining capacitance would be Co/2
Now both of these capacitance would be in Parallel therefore
Net capacitance
= Co/2 + kCo /2 = (Co /2 )(1+k)